# Rate-reliability functions

### v2

@@ Abstract
We investigate deterministic identification over arbitrary memoryless channels under the constraint that the error probabilities of first and second kind are exponentially small in the block length $n$, controlled by reliability exponents $E_{1},E_{2}\geq 0$. In contrast to the regime of slowly vanishing errors, where the identifiable message length scales as $\Theta(n\log n)$, here we find that for positive exponents linear scaling is restored, now with a rate that is a function of the reliability exponents. We give upper and lower bounds on the ensuing rate-reliability function in terms of (the logarithm of) the packing and covering numbers of the channel output set, which for small error exponents $E_{1},E_{2}>0$ can be expanded in leading order as the product of the Minkowski dimension of a certain parametrisation the channel output set and $\log\min\{E_{1},E_{2}\}$. These allow us to recover the previously observed slightly superlinear identification rates, and offer a different perspective for understanding them in more traditional information theory terms. We further illustrate our results with a discussion of the case of dimension zero, and extend them to classical-quantum channels and quantum channels with tensor product input restriction.
@@
Shannon theory; identification via channels; finite block length; memoryless systems; quantum information.
@@ I Introduction
In Shannon’s fundamental model of communication [Shannon:TheoryCommunication], the receiver aims to faithfully recover an original message sent through $n$ uses of a memoryless noisy channel $W:\mathcal{X}\rightarrow\mathcal{Y}$, given by a probability transition kernel between measurable spaces $\mathcal{X}$ and $\mathcal{Y}$. The maximum number $M$ of transmittable messages scales exponentially with $n$, and thus we define the (linear-scale) rate of a transmission code as $R=\frac{1}{n}\log M$, with $\log$ the binary logarithm. The maximum rate for asymptotically faithful transmission – that is, with arbitrarily small probability of error for sufficiently long code words – is known as the capacity $C$ of the channel [Shannon:TheoryCommunication, Wolfowitz:converse].
This model contrasts with the identification task in which instead of recovering the initial message, the receiver is only interested in knowing whether it is equal to one particular message of their choice or not. Jaja demonstrated in [Ja:ID_easier] that this task involves lower communication complexity than Shannon’s original transmission scenario. Building on this result, Ahlswede and Dueck [AD:ID_ViaChannels] fully characterized the identification problem and analyzed its performance over general discrete memoryless channels (DMCs) using a randomized encoder. They showed that identification codes can achieve doubly exponential growth of the number $N$ of messages as a function of the block length $n$, i.e. $\log N\sim 2^{nR}$. In other words, we can identify exponentially more messages than we can transmit. As this result relies crucially on randomized encoding, it is natural to ask what happens if we impose deterministic encoding.
@@
$\forall j\neq k\in[N]\quad W_{u_{j}}(\mathcal{E}_{j})\geq 1-\lambda_{1},\quad W_{u_{j}}(\mathcal{E}_{k})\leq\lambda_{2},$
@@
It was initially observed that deterministic identification leads to much poorer results in terms of code size scaling in the block length [AD:ID_ViaChannels, AC:DI] than the unrestricted randomized identification. Indeed, DI over discrete memoryless channels can only lead to linear scaling of the message length, $\log N\sim Rn$ [SPBD:DI_power] as in Shannon’s communication paradigm (albeit with a higher rate $R$). Despite this poorer performance, interest in deterministic codes has recently been renewed, as they have proven to be easier to implement and simulate [DI_simpler_impl], to explicitly construct [DI_explicit_construction], and offer reliable single-block performance [AD:ID_ViaChannels]; see in particular [VDTB:practical-DI] for identification codes. Also, surprisingly, certain channels with continuous input alphabets have DI codes governed by a slightly superlinear scaling in the block length: $\log N\sim Rn\log n$. This was first observed for Gaussian channels [SPBD:DI_power], Gaussian channels with both fast and slow fading [DI-fading, VDB:DI-fading-new], and Poisson channels [DI-poisson, DI-poisson_mc]. This behaviour differs from that of randomized identification, where the scaling remains the same as for DMCs [Han:book, LDB:mc].
In our recent work [CDBW:DI_classical, CDBW:DI_proceedings] we have generalized all these particular DI results, proving that the superlinear scaling observed in these special cases is actually a feature exhibited by general channels with discrete output. For the rest of the paper we will in fact assume that $\mathcal{Y}$ is finite. The corresponding superlinear pessimistic and optimistic capacities (see [Ahlswede2006] or the discussion in [CDBW:DI_classical, Sect. 7.1] for further insight on this topic) can be bounded respectively in terms of the lower and upper covering dimension (an introduction and definitions of these concepts can be found in Section II), $\underline{d}_{M}$ and $\overline{d}_{M}$, of a certain algebraic transformation of the set of output probability distributions of the channel:
@@
When $d=0$, like in the case of the discrete memoryless channel, a more general and abstract theorem is provided to meaningfully bound the size of the code at the correct (slower) scaling [CDBW:DI_classical, Thm. 5.11].
While the DI capacity is defined as the maximum achievable (superlinear) rate for reliable communication – where the error probability can be made arbitrarily small with sufficiently large $n$ – it does a priori not provide information on how quickly the errors vanish. In contrast, it is well-known in the communication setting [Shannon:TheoryCommunication, Gallager:ReliabilityBook] that at (linear scale) rates below capacity, $0\leq R<C$, the error probability decreases exponentially with the block length: $\lambda\sim 2^{-En}$. In the present paper we initiate the study of the performance of DI codes when $\lambda_{1}$ and $\lambda_{2}$ are exponentially small, in an attempt to better understand the superlinear rates at slowly vanishing errors. As it turns out, we can rely on many ideas and tools from our previous work [CDBW:DI_classical], as it was written with an eye to finite block length. The present work extends the results from [CDBW:DI_classical], which are recovered from here in the limit of error exponents slowly going to 00.
Anticipating our subsequent results, we associate to every $(n,N,\lambda_{1},\lambda_{2})$-DI code the triple $(R(n),E_{1}(n),E_{2}(n))$ of linear-scale message and error rates:
$R(n):=\frac{1}{n}\log N,\,\,E_{1}(n):=\frac{1}{n}\log\frac{1}{\lambda_{1}},\,\,E_{2}(n):=\frac{1}{n}\log\frac{1}{\lambda_{2}}.$
We call $(R,E_{1},E_{2})$ an (asymptotically) achievable triple if there exist deterministic identification codes with $(R(n),E_{1}(n),E_{2}(n))$ converging to $(R,E_{1},E_{2})$ for $n\rightarrow\infty$. The set of all achievable triples will be referred to as the achievable region $\mathcal{R}(W):=\{(R,E_{1},E_{2})\text{ achievable}\}$. The rate-reliability function is then
$R(E_{1},E_{2}):=\sup_{E_{1},E_{2}}R\quad\text{s.t.}\quad(R,E_{1},E_{2})\in\mathcal{R}(W).$
Similarly, we can define an optimistic achievable region ¯$\overline{\mathcal{R}}(W)\supseteq\mathcal{R}(W)$ by allowing the convergence to happen only for a subsequence $n_{k}\rightarrow\infty$ of block lengths. That is to say, ¯$\overline{\mathcal{R}}(W)$ is the set of all triples $(R,E_{1},E_{2})$ such that there exist codes of block length $n_{k}\rightarrow\infty$ with $(R{(n_{k})},E_{1}{(n_{k})},E_{2}{(n_{k})})\rightarrow(R,E_{1},E_{2})$. The optimistic rate-reliability function is thus
$\overline{R}(E_{1},E_{2}):=\sup_{E_{1},E_{2}}R\quad\text{s.t.}\quad(R,E_{1},E_{2})\in\overline{\mathcal{R}}(W).$
@@
Figure 1: Schematic of the rate-reliability function for transmission and DI. The rate-reliability function for transmission (blue line) is a monotonically non-increasing function with different regions depending on the optimum coding strategies (see the tendency change from linear to quadratic behaviour marked by the grey dotted line). For error exponents going to zero (very slow error decrease with the block length $n$), the linear rate converges to the transmission capacity $C(W)$; no higher rates are possible. Here, we provide bounds on the rate-reliability function for DI (red line) that show how the rates go to infinity logarithmically in the limit of error exponents $E\rightarrow 0$, corresponding to the superlinearity of the rates in that regime. Note that the DI curve must dominate the transmission curve everywhere, since an $(n,N,\lambda)$-transmission code is automatically an $(n,N,\lambda,\lambda)$-DI code.
@@
In the rest of the paper, after giving the necessary technical tools in Section II, we present our rate-reliability bounds in Section III. We remark already here that while those bounds are stated for general exponents, they are most suitable for exponents below a certain threshold. After that, in Section IV we discuss the bounds from three different angles: in Subsection IV-A we show that for all (sufficiently small) $E_{1},E_{2}>0$, the rate-reliability functions $R(E_{1},E_{2})$ and $\overline{R}(E_{1},E_{2})$ are bounded away from 00 and $\infty$; and by letting $E_{1}(n),E_{2}(n)$ vanish slowly, we recover the main superlinear capacity results from [CDBW:DI_classical] (see the schematic Figure 1). In Subsection IV-B we show how to deal with channels characterised by dimension zero, where the previous bounds seem not to give the full picture of what is happening. In Subsection IV-D we discuss the regime of large error exponents (quickly vanishing errors). Finally, we generalize the bounds to the quantum setting for classical-quantum channels, and general quantum channels under the restriction that only product states are used on the input (Subsection IV-C). We close with a brief discussion and conclusions in Section V.
@@ II Preliminaries
Our technical results rely on the methods developed in [CDBW:DI_classical]. These are mainly based on metric studies of the output probability sets. The essential distance measures needed are defined next. The total variation distance is a statistical distance measure which coincides with half the $L^{1}$ distance between the probability mass functions. Let $P$ and $Q$ be two probability distributions defined on a finite or countably infinite measurable space $\mathcal{L}$, then the total variation distance is defined as
$\frac{1}{2}\|P-Q\|_{1}=\sum_{\ell\in\mathcal{L}}\frac{1}{2}|P(\ell)-Q(\ell)|.$
@@
Finally, following [CDBW:DI_classical, Sect. 3] we use a metric on a modified output set that comes out naturally for product distributions, and will allow us to work in a Euclidean space (enabling the use of the Minkowski dimensions defined above). For a probability distribution $P\in\mathcal{P}(\mathcal{L})$, define the unit vector $\sqrt{P}:=(\sqrt{P(\ell)}:\ell\in\mathcal{L})\in\mathbb{R}^{\mathcal{L}}$. The image of $\mathcal{P}(\mathcal{L})$ under the square root map is the non-negative orthant sector of the unit hypersphere in $\mathbb{R}^{|\mathcal{L}|}$, which we shall denote $S_{+}(|\mathcal{L}|,1)$. This obeys the relation
@@
For a channel $W:\mathcal{X}\rightarrow\mathcal{Y}$, we denote the output probability set $\widetilde{\mathcal{X}}:=W(\mathcal{X})\subset\mathcal{P}(\mathcal{Y})$, and hence have
$\sqrt{\!\widetilde{\mathcal{X}}}=\left\{\sqrt{W_{x}}:x\in\mathcal{X}\right\}\subset S_{+}(|\mathcal{Y}|,1).$
@@
The nice properties exhibited by typical sets (they collect almost all the probability, which is almost evenly distributed among their elements, cf. [winter:PhDThesis, Lemmas I.11 & I.12]) allow us to bound the errors of first and second kinds (when choosing $\mathcal{T}_{u_{j}}^{\delta}$ as decision rule) through the following lemmas, which will be instrumental in the next section.
Lemma 2 ([CDBW:DI_classical, Lemma 2.1]).
For an arbitrary channel $W:\mathcal{X}\rightarrow\mathcal{Y}$, arbitrary block length $n$, $0<\delta\leq\sqrt{n}\log|\mathcal{Y}|$, and for any $x^{n}\in\mathcal{X}^{n}$, we have
$W_{x^{n}}(\mathcal{T}_{x^{n}}^{\delta})\geq 1-2\exp\left(-\delta^{2}/36K(|\mathcal{Y}|)\right),$
where $K(d)=(\log\max\{d,3\})^{2}$.
Lemma 3 ([CDBW:DI_classical, Lemma 3.1]).
@@
$\begin{split}W_{{x^{\prime}}^{n}}(\mathcal{T}_{x^{n}}^{\delta})&\leq 2\exp\left(-\delta^{2}/36K(|\mathcal{Y}|)\right)\\ &\phantom{\leq.}+\epsilon\left(1+2^{2\delta\sqrt{n}}2^{H(W_{x^{n}})-H(W_{{x^{\prime}}^{n}})}\right).\end{split}$
For both the coding and converse part, we will need the packing and covering of general sets in arbitrary dimensions. For a non-empty bounded subset $F$, the covering problem consists in finding the minimum number $\Gamma_{\delta}(F)$ of closed balls of radius $\delta$ centered at points in $F$ such that their union contains $F$; in contrast, the packing problem consists in finding the maximum number $\Pi_{\delta}(F)$ of pairwise disjoint open balls of radius $\delta$ centered at points in $F$. These packing and covering numbers are fundamental in geometry as they can be used to define the Minkowski dimension (also known as covering, Kolmogorov, or entropy dimension) of a subset in Euclidean space as
@@
While for smooth manifolds the topological and Minkowski dimensions coincide, the latter is especially relevant in theory of fractals, as it captures the geometric complexity of the set beyond the scope of the classical topological dimension, see the excellent textbook [Falconer:fractal], as well as [Robinson:dimensions, Fraser:dimensions] for further insight. For even more general and irregular sets, it is quite common that the above limit [Eq. (6)] does not exist. In that case, we define the upper and lower Minkowski dimensions through the limit superior and limit inferior, respectively:
@@
The dimensions of the image of output probability set $\widetilde{\mathcal{X}}$ and its square root $\sqrt{\!\widetilde{\mathcal{X}}}$ are subject to the inequalities $d_{M}(\widetilde{\mathcal{X}})\leq d_{M}\!\left(\!\sqrt{\!\widetilde{\mathcal{X}}}\right)\leq 2d_{M}(\widetilde{\mathcal{X}})$. This is because $\widetilde{\mathcal{X}}$ is a Lipschitz-continuous image of $\sqrt{\!\widetilde{\mathcal{X}}}$; and conversely, $\sqrt{\!\widetilde{\mathcal{X}}}$ is a $\frac{1}{2}$-Hölder-continuous image of $\widetilde{\mathcal{X}}$, cf. [Robinson:dimensions, Fraser:dimensions].

### v4

# Rate-reliability tradeoff

@@ Abstract
We investigate deterministic identification over arbitrary memoryless channels under the constraint that the error probabilities of first and second kind are exponentially small in the block length $\mathbf{n}$, controlled by reliability exponents $\mathbf{E_{1},E_{2}\geq 0}$. In contrast to the regime of slowly vanishing errors, where the identifiable message length scales linearithmically as $\mathbf{\Theta(n\log n)}$, here we find that for positive exponents linear scaling is restored, now with a rate that is a function of the reliability exponents. We give upper and lower bounds on the ensuing rate-reliability function in terms of (the logarithm of) the packing and covering numbers of the channel output set, which for small error exponents $\mathbf{E_{1},E_{2}>0}$ can be expanded in leading order as the product of the Minkowski dimension of a certain parametrisation the channel output set and $\mathbf{\log\min\{E_{1},E_{2}\}}$. These allow us to recover the previously observed slightly superlinear identification rates, and offer a different perspective for understanding them in more traditional information theory terms. We also show that even if only one of the two errors is required to be exponentially small, the linearithmic scaling is lost. We further illustrate our results with a discussion of the case of dimension zero, and extend them to classical-quantum channels and quantum channels with tensor product input restriction.
@@
Shannon theory; identification via channels; rate-reliability function; finite block length; memoryless systems; quantum information.
@@ I Introduction
I-A Background and previous work
In Shannon’s fundamental model of communication [2], the receiver aims to faithfully recover an original message sent through $n$ uses of a memoryless noisy channel $W:\mathcal{X}\rightarrow\mathcal{Y}$, given by a probability transition kernel between measurable spaces $\mathcal{X}$ and $\mathcal{Y}$. The maximum number $M$ of transmittable messages scales exponentially with $n$, and thus we define the (linear-scale) rate of a transmission code as $R=\frac{1}{n}\log M$, with $\log$ the binary logarithm. The maximum rate for asymptotically faithful transmission – that is, with arbitrarily small probability of error for sufficiently long code words – is known as the capacity $C$ of the channel [2, 3].
This model contrasts with the identification task in which instead of recovering the initial message, the receiver is only interested in knowing whether it is equal to one particular message of their choice or not. Jaja demonstrated in [4] that this task involves lower communication complexity than Shannon’s original transmission scenario. Building on this result, Ahlswede and Dueck [5] fully characterized the identification problem and analyzed its performance over general discrete memoryless channels (DMCs) using a randomized encoder. They showed that identification codes can achieve doubly exponential growth of the number $N$ of messages as a function of the block length $n$, i.e. $\log N\sim 2^{nR}$. In other words, we can identify exponentially more messages than we can transmit. In general, randomized identification codes can achieve this exponential scaling [6, 7]. As these results rely crucially on randomized encodings, it is natural to ask what happens if we impose a deterministic approach.
@@ I-A Background and previous work
$\forall j\neq k\in[N],\quad W_{u_{j}}(\mathcal{E}_{j})\geq 1-\lambda_{1},\quad W_{u_{j}}(\mathcal{E}_{k})\leq\lambda_{2},$ (1)
@@
Notice the structural differences between transmission and identification codes: in transmission, there is a single type of error (incorrect message decoding) while in identification, two distinct errors arise. A missed identification $\lambda_{1}$ occurs when the correct message is not identified, and a wrong identification $\lambda_{2}$ when an incorrect message is mistakenly identified. The latter is directly tied to the fact that identification decoding sets $\mathcal{E}_{j}$ may overlap, unlike the decoding regions required in transmission which are disjoint.
It was initially observed that DI leads to much poorer results in terms of code size scaling in the block length [5, 8] than the unrestricted randomized identification. Indeed, DI over discrete memoryless channels can only lead to linear scaling of the message length, $\log N\sim Rn$ [9] as in Shannon’s communication paradigm (albeit with a higher rate $R$). Despite this poorer performance, interest in deterministic codes has recently been renewed, as they have proven to be easier to implement and simulate [10], to explicitly construct [11], and offer reliable single-block performance [5]; see in particular [12] for identification codes. Also, surprisingly, certain channels with continuous input alphabets have DI codes governed by a slightly superlinear scaling in the block length: $\log N\sim Rn\log n$ (commonly referred to as linearithmic). This was first observed for Gaussian channels [9], Gaussian channels with both fast and slow fading [13, 14], and Poisson channels [15, 16]. In our recent work [17, 18] we have generalized all these particular DI results, proving that the linearithmic scaling observed in these special cases is actually a feature exhibited by general channels with discrete output. For the rest of the paper we will in fact assume that $\mathcal{Y}$ is finite.
It is important to emphasize the fundamental implications of different scalings beyond the value of the rate expressions. When a code achieves superlinear scaling – for example linearithmic $\log N(n)\sim Rn\log n$ –, the number of identifiable messages grows asymptotically faster than in the linear regime $\log N(n)\sim R^{\prime}n$, regardless of how large the linear rate constant might be. That is, even if a linear code achieves a higher rate $R^{\prime}>R$, the superlinear code will eventually yield infinitely more identifiable messages as $n\to\infty$. This distinction justifies a careful attention to the growth regime. Specifically when calculating the capacity of a channel, it is crucial to define it in the correct scaling. Indeed, if we have a linearithmic growth of code length, the linear (exponential) scale rate will trivially be $\infty$ (0). To define a meaningful capacity value, it is essential to choose the suitable scale – in the present case linearithmic.
In [17], linearithmic DI capacities $\dot{C}_{\text{DI}}$ of general channels were analysed in the pessimistic and optimistic settings; the formal definition of the capacities can be found in Subsection IV-A (for further insight on these settings see [19] or the discussion in [17, Sect. 7.1]). It was shown that they can be respectively bounded in terms of the lower and upper covering dimension $\underline{d}_{M}$ and $\overline{d}_{M}$ [an introduction and the definitions of these concepts is found in Section II below, see Equations (2) and (3)] of a certain algebraic transformation of the set of output probability distributions of the channel:
@@
When $d=0$, such as in the case of the discrete memoryless channel, a more general and abstract theorem is provided to meaningfully bound the size of the code at the correct (in general slower, for DMCs in fact linear) scaling [17, Thm. 5.11].
I-B Motivation and preliminary problem setting
While the DI capacity is defined as the maximum achievable (linearithmic) rate for reliable identification – where the error probability can be made arbitrarily small with sufficiently large $n$ –, it does a priori not provide information on how quickly the errors vanish. In contrast, it is well-known in the communication setting [2, 20] that at (linear-scale) rates below capacity, $0\leq R<C$, the error probability decreases exponentially with the block length: $\lambda\sim 2^{-En}$. In the present paper we initiate the study of the performance of DI codes when $\lambda_{1}$ and/or $\lambda_{2}$ are exponentially small, in an attempt to better understand the nature of the superlinear rates.
Anticipating our subsequent results, we associate to every $(n,N,\lambda_{1},\lambda_{2})$-DI code (recall Definition 1) the triple of linear-scale message and error rates $(R(n),E_{1}(n),E_{2}(n))$, where:
$\!\!\!\!R(n)\!:=\!\frac{1}{n}\log N,\,\,E_{1}(n)\!:=\!\frac{1}{n}\log\frac{1}{\lambda_{1}},\,E_{2}(n)\!:=\!\frac{1}{n}\log\frac{1}{\lambda_{2}}.$ (4)
The main object of study in this work is the tradeoff between the linear-scale rate functions $R(n)$ and the error exponents $E_{1,2}(n)$ (with the notation $E_{1,2}$ meaning that the same equation holds for both $E_{1}$ and $E_{2}$), first in the finite block length regime, and then on the asymptotic limit $n\rightarrow\infty$. The error exponents can of course be functions of $n$ as long as $\lim_{n\rightarrow\infty}nE_{1,2}(n)=\infty$, since we need the errors to vanish in the asymptotic limit; in other words, $E_{1,2}(n)\geq\omega(1/n)$.
We call $(R,E_{1},E_{2})$ an (asymptotically) achievable triple if there exist DI codes with $(R(n),E_{1}(n),E_{2}(n))$ converging to $(R,E_{1},E_{2})$ for $n\rightarrow\infty$. The set of all achievable triples will be referred to as the achievable region $\mathcal{R}(W):=\{(R,E_{1},E_{2})\text{ achievable}\}$. The rate-reliability function is then
$R(E_{1},E_{2}):=\sup_{E_{1},E_{2}}R\quad\text{s.t.}\quad(R,E_{1},E_{2})\in\mathcal{R}(W).$ (5)
Similarly, we can define an optimistic achievable region ¯$\overline{\mathcal{R}}(W)\supseteq\mathcal{R}(W)$ by allowing the convergence to happen only for a subsequence $n_{k}\rightarrow\infty$ of block lengths. That is to say, ¯$\overline{\mathcal{R}}(W)$ is the set of all triples $(R,E_{1},E_{2})$ such that there exist deterministic identification codes of block length $n_{k}\rightarrow\infty$ with $(R{(n_{k})},E_{1}{(n_{k})},E_{2}{(n_{k})})\rightarrow(R,E_{1},E_{2})$. The optimistic rate-reliability function is thus
$\overline{R}(E_{1},E_{2}):=\sup_{E_{1},E_{2}}R\quad\text{s.t.}\quad(R,E_{1},E_{2})\in\overline{\mathcal{R}}(W).$ (6)
@@ I-B Motivation and preliminary problem setting
While in our previous work [17] we had no restrictions on the identification errors other than them being sufficiently small as $n\rightarrow\infty$, here we study in depth the performance of codes with different error settings. We start with the case where both errors vanish exponentially, for which we provide lower and upper bounds on the trade-off between rate and error exponents in the finite block length regime. The study of these bounds in the asymptotic limit allow us to show not only that codes with exponentially vanishing errors can only have linear scaling, but also to find the error conditions under which the main superlinear results in [17] can be recovered. We furthermore study the cases where only one of the two errors is constrained to vanish exponentially, showing that linearithmic scaling is also lost. Despite the current analysis being much more complete both in terms of error settings and non-asymptotic regimes, we can rely on many ideas and tools from our previous work [17], as it was written with an eye to finite block length.
@@
Figure 1: Schematic of the rate-reliability function for transmission and DI. The rate-reliability function for transmission (blue line) is monotonically non-increasing with different regions depending on the optimum coding strategies (see the tendency change from linear to quadratic behaviour marked by the grey dotted line). For error exponents going to zero (very slow error decrease with the block length $n$), the linear rate converges to the transmission capacity $C(W)$; no higher rates are possible. Here, in Sections III and IV, we provide bounds on the rate-reliability function for DI (red line) that show how the rate (defined in the linear scale) goes to infinity logarithmically in the error exponents $E_{1}=E_{2}=E\approx 0$, corresponding to the linearithmic code length in that regime. Note that the DI curve must dominate the transmission curve everywhere, since an $(n,N,\lambda)$-transmission code is automatically an $(n,N,\lambda,\lambda)$-DI code – indeed, if we can decode a message, we can identify it.
I-C Outline
After giving the necessary technical tools in Section II, we present upper and lower bounds on the non-asymptotic rates as functions of the error exponents in Section III. We remark already here that while these bounds are stated for general exponents, they are most suitable for exponents below a certain threshold. A full account of this fact and alternative analyses for the regime of large error exponents are given in the subsequent sections.
@@ I-C Outline
In Section IV, we discuss the bounds from three different angles: in IV-A we show that for all (sufficiently small) $E_{1},E_{2}>0$, the rate-reliability functions $R(E_{1},E_{2})$ and $\overline{R}(E_{1},E_{2})$ are bounded away from 0 and $\infty$; and by letting $E_{1}(n)$ and $E_{2}(n)$ vanish slowly, we recover the main linearithmic capacity results from [17] (see the schematic Figure 1). In Subsection IV-B we show how to deal with channels characterised by dimension zero, where the previous bounds seem not to give the full picture of what is happening. Finally, in IV-C we discuss the regime of large error exponents (quickly vanishing errors) and why the bounds in III are not suitable there.
Section V studies the case of asymmetric error exponents, that is, when we constrain one of the errors to vanish exponentially while the other is just bounded. While the results are used to broaden the previous discussion, we need a completely different approach, hence the separate section.
Finally, in Section VI, we generalize the main bounds in Section III to the quantum setting for classical-quantum channels, and general quantum channels under the restriction that only product states are used on the input. We close with a brief discussion and open problems in Section VII.
@@ II Preliminaries
As we have already mentioned, a number of technical ideas in this work rely on the methods developed in [17]. These are mainly based on metric studies of the output probability sets. The essential distance measures needed are defined next: the total variation distance is a statistical distance measure which coincides with half the $L^{1}$ distance between the probability mass functions. Let $P$ and $Q$ be two probability distributions defined on a finite or countably infinite measurable space $\mathcal{L}$, then the total variation distance is defined as
$\!\!\!\frac{1}{2}\|P-Q\|_{1}\!:=\sup_{L\in\mathcal{L}}|P(L)-Q(L)|\!=\!\!\sum_{\ell\in\mathcal{L}}\frac{1}{2}|P(\ell)-Q(\ell)|.$ (7)
@@
Finally, following [17, Sect. 3] we use a metric on a modified output set that comes out naturally for product distributions, and will allow us to work in a Euclidean space (enabling the use of the Minkowski dimensions which already appeared in Subsection I-A above, and are defined below). For a probability distribution $P\in\mathcal{P}(\mathcal{L})$, define the unit vector $\sqrt{P}:=(\sqrt{P(\ell)}:\ell\in\mathcal{L})\in\mathbb{R}^{\mathcal{L}}$. The image of $\mathcal{P}(\mathcal{L})$ under the square root map is the non-negative orthant sector of the unit hypersphere in $\mathbb{R}^{|\mathcal{L}|}$, which we shall denote $S_{+}(|\mathcal{L}|,1)$. As proven in [17], these objects obey the relation
@@
For a channel $W:\mathcal{X}\rightarrow\mathcal{Y}$, we denote the output probability set $\widetilde{\mathcal{X}}:=W(\mathcal{X})\subset\mathcal{P}(\mathcal{Y})$, with $W_{x}$ the output distribution conditioned on an input $x$ as described in Definition 1. Hence, we have
$\sqrt{\!\widetilde{\mathcal{X}}}=\left\{\sqrt{W_{x}}:x\in\mathcal{X}\right\}\subset S_{+}(|\mathcal{Y}|,1).$ (10)
@@
The nice properties exhibited by typical sets (they collect almost all the probability, which is almost evenly distributed among their elements, cf. [21, Lemmas I.11 & I.12]) allow us to bound the errors of first and second kinds (when choosing $\mathcal{T}_{u_{j}}^{\delta}$ as decision rule) through the following lemmas, which will be instrumental in the next section.
Lemma 2 ([17, Lemma 2.1]).
Let $K(d):=(\log\max\{d,3\})^{2}$. Then, for an arbitrary channel $W:\mathcal{X}\rightarrow\mathcal{Y}$, arbitrary block length $n$, $0<\delta\leq\sqrt{n}\log|\mathcal{Y}|$, and for any $x^{n}\in\mathcal{X}^{n}$, we have
$W_{x^{n}}(\mathcal{T}_{x^{n}}^{\delta})\geq 1-2\exp\left(-\delta^{2}/36K(|\mathcal{Y}|)\right).$ (12)
Lemma 3 ([17, Lemma 3.1]).
@@
$\begin{split}W_{{x^{\prime}}^{n}}(\mathcal{T}_{x^{n}}^{\delta})&\leq 2\exp\left(-\delta^{2}/36K(|\mathcal{Y}|)\right)\\ &\phantom{\leq.}+\epsilon\left(1+2^{2\delta\sqrt{n}}2^{H(W_{x^{n}})-H(W_{{x^{\prime}}^{n}})}\right).\end{split}$ (13)
For both the coding and converse part, we will need the packing and covering of general sets in arbitrary dimensions. For a non-empty bounded subset $F$, the covering problem consists in finding the minimum number $\Gamma_{\delta}(F)$ of closed balls of radius $\delta$ centered at points in $F$ such that their union contains $F$; in contrast, the packing problem consists in finding the maximum number $\Pi_{\delta}(F)$ of pairwise disjoint open balls of radius $\delta$ centered at points in $F$. The packing and covering numbers are not equal in general, but they exhibit the same asymptotic behaviour in the exponent as $\delta\to 0$ [22]. These are fundamental objects in geometry as they can be used to define the Minkowski dimension (a.k.a covering, Kolmogorov, or entropy dimension) of a subset in Euclidean space as
@@
While for smooth manifolds the topological and Minkowski dimensions coincide, the latter is especially relevant in theory of fractals, as it captures the geometric complexity of the set beyond the scope of the classical topological dimension, see the excellent textbook [22], as well as [23, 24] for further insight. For even more general and irregular sets, it is quite common that the above limit [Eq. (14)] does not exist. In that case, we define the upper and lower Minkowski dimensions through the limit superior and limit inferior, respectively:
@@
The dimensions of the image of output probability set $\widetilde{\mathcal{X}}$ and its square root $\sqrt{\!\widetilde{\mathcal{X}}}$ are subject to the inequalities $d_{M}(\widetilde{\mathcal{X}})\leq d_{M}\!\left(\!\sqrt{\!\widetilde{\mathcal{X}}}\right)\leq 2d_{M}(\widetilde{\mathcal{X}})$. This is because $\widetilde{\mathcal{X}}$ is a Lipschitz-continuous image of $\sqrt{\!\widetilde{\mathcal{X}}}$; and conversely, $\sqrt{\!\widetilde{\mathcal{X}}}$ is a $\frac{1}{2}$-Hölder-continuous image of $\widetilde{\mathcal{X}}$, cf. [23, 24].